# Optimizing Learning for Robust Hyperbolic Deep Learning in Computer Vision

### v2

@@ Abstract
Hyperbolic deep learning has become a growing research direction in computer vision for the unique properties afforded by the alternate embedding space. The negative curvature and exponentially growing distance metric provide a natural framework for capturing hierarchical relationships between datapoints and allowing for finer separability between their embeddings. However, these methods are still computationally expensive and prone to instability, especially when attempting to learn the negative curvature that best suits the task and the data. Current Riemannian optimizers do not account for changes in the manifold which greatly harms performance and forces lower learning rates to minimize projection errors. Our paper focuses on curvature learning by introducing an improved schema for popular learning algorithms and providing a novel normalization approach to constrain embeddings within the variable representative radius of the manifold. Additionally, we introduce a novel formulation for Riemannian AdamW, and alternative hybrid encoder techniques and foundational formulations for current convolutional hyperbolic operations, greatly reducing the computational penalty of the hyperbolic embedding space. Our approach demonstrates consistent performance improvements across both direct classification and hierarchical metric learning tasks while allowing for larger hyperbolic models.
@@ 1 Introduction
With the recent rise in the use of hyperbolic manifolds for deep representation learning, there is a growing need for efficient, flexible components that can fully exploit these spaces without sacrificing stability. This has led researchers to focus on two main derivations of hyperbolic space: the Poincaré manifold and the hyperboloid. The Poincaré ball, equipped with a gyrovector space, supports various well-defined operations, including generalized vector addition and multiplication, but it suffers from significant stability issues. On the other hand, the hyperboloid, or Lorentz space, lacks these operations but offers much better operation stability, as demonstrated in the study by Mishne et al. (2022).
To address this gap, previous works have sought to provide Lorentzian definitions for common deep learning operations such as the feed-forward layer (Chen et al., 2022; Dai et al., 2021; Ganea et al., 2018), convolutional layer (Chen et al., 2022; Qu & Zou, 2023; Dai et al., 2021), and MLRs (Bdeir et al., 2024). This increased focus on hyperbolic modeling has led to its gradual integration into computer vision architectures, as detailed in the survey by Mettes et al. (2023). Specifically, the hyperboloid model has been employed as a sampling space for VAEs (Nagano et al., 2019), a decoder space for vision tasks in hybrid settings (Guo et al., 2022; Liu et al., 2020; Khrulkov et al., 2020; Qu & Zou, 2022), and ultimately for fully hyperbolic Lorentzian vision encoders (Bdeir et al., 2024) simultaneously with its Poincaré counterpart (van Spengler et al., 2023).
This paper furthers the development of hyperbolic learning for vision tasks, specifically for the Lorentz manifold. Our primary focus is on the challenge of learning the manifold’s negative curvature. The driving principle behind this, is that the model embeddings may exhibit varying degrees of hyperbolicity depending on the innate hierarchies in the datapoints themselves, the problem task that is being considered, and the specific locations of hyperbolic operation integrations. To accomodate for this, we can adjust the embedding space’s hyperbolic metric to be less or more Euclidean which accounts for the modeling requirements. We also build on the idea of separate manifolds for separate main blocks in the architecture further increasing representative flexibility.
We also recognize that despite recent advances, Lorentz models continue to struggle with issues of high computational costs. We attempt to isolate and alleviate the main factors leading to numerical inaccuracies and computational overhead overall, and more particularly when modeling data in higher-dimensional embedding spaces and when learning the curvatures. Our contributions can then be summed up as:
@@
    We propose a formulation for Riemannian AdamW and an alternative schema for Riemannian optimizers that accounts for manifold curvature learning.
@@
    We provide a more efficient convolutional layer approach that is able to leverage the highly optimized existing implementations.
@@
    We empirically show the effectiveness of combining these approaches using classical image classification tasks and hierarchical metric learning problems.
@@ 2 Related Work
Hyperbolic Embeddings in Computer Vision
@@
Metric Learning
Metric learning relies on the concept of structuring the distribution in the embedding space so that related data points are positioned closer together, while less related points are placed further apart. To facilitate this process, numerous studies have introduced additional loss functions that explicitly encourage this behavior. Contrastive losses, for instance, operate on pairs of data points and propose a penalty that is proportional to the distances between negative pairs and inversely proportional to the distance between positive pairs (Chopra et al., 2005). Triplet loss extends this idea by considering sets of three points: an anchor, a positive sample, and a negative sample (Wang et al., 2014). Instead of changing the distances between points absolutely, it ensures that the distance between the anchor and the positive sample is less than the distance between the anchor and the negative sample, plus a margin, thus enforcing a relational criterion.
These approaches have also been adapted to hierarchical problem settings under hyperbolic manifolds (Yang et al., 2022; Kim et al., 2023). Notably, Kim et al. (2023) developed a method for learning continuous hierarchical representations using a deep learning, data-mining-like approach that relies on the innate relationships of the embeddings rather than their labels. They employ a proxy-based method that models the data on the Poincaré ball, facilitating a more natural extension to hierarchical tasks. Building on this, we extend the approach by modeling the loss function in the Lorentz manifold and incorporating a learnable curvature to better handle data with varying levels of hierarchy.
@@ 3 Methodology
Hyperbolic embeddings offer significant advantages for computer vision tasks that benefit from hierarchical structure modeling, but their effectiveness is hindered by practical challenges, including mathematical inconsistencies in curvature learning, the absence of a Riemannian AdamW optimizer, stability issues with float32 precision, and high computational costs from inefficient convolutions.
To address these issues, the following sections introduce four key advancements: a more stable curvature learning method, a Riemannian AdamW optimizer, a normalization scheme for improved float32 precision stability, and efficient hyperbolic convolutions to reduce memory and computational costs.
3.1 Background
The hyperbolic space is a Riemannian manifold with a constant negative sectional curvature $c<0$. There are many conformal models of hyperbolic space but we focus our work on the hyperboloid, or Lorentz manifold. The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}\in\mathbb{R}^{n+1}~{}|~{}\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=\frac{-1}{K},~{}x_{t}>0\}$ where $\frac{-1}{K}=c$, and with the Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\mathrm{diag}(-1,1,\dots,1)$. This models the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}=[\sqrt{K},0,\cdots,0]^{T}$. We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector, the space dimension $\bm{x_{s}}$. The Lorentzian inner product then becomes $\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}=\bm{x}^{T}\mathrm{diag}(-1,1,\cdots,1)\bm{y}.$ We now define the common hyperbolic operations in the Lorentz space.
@@ 3.1 Background
Distance in hyperbolic space is the magnitude of the geodesic forming the shortest path between two points. Let $\bm{x},\bm{y}\in\mathbb{L}^{n}_{K}$, the distance between them is given by $d_{\mathbb{L}}(\bm{x},\bm{y})=\sqrt{K}\cosh^{-1}(\frac{-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}{K})$. We also define the square distance by Law et al. (2019) as $d^{2}_{\mathbb{L}}(\bm{x},\bm{y})=||\bm{x}-\bm{y}||_{\mathbb{L}}^{2}=-2K-2\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$.
@@
Seeing as the Lorentz space is a Riemannian manifold, it is locally Euclidean. This can best be described through the tangent space $\mathbb{T}_{\bm{x}}\mathcal{M}$, a first-order approximation of the manifold at a given point $\bm{x}$. The exponential map, $\exp^{K}_{\bm{x}}\left(\bm{z}\right):\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}\rightarrow\mathbb{L}^{n}_{K}$ is then the operation that maps a tangent vector $\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}$ onto the manifold through $\exp_{\bm{x}}^{K}(\bm{z})=\cosh(\alpha)\bm{x}+\sinh(\alpha)\frac{\bm{z}}{\alpha},\text{ with }\alpha=\sqrt{1/K}||\bm{z}||_{\mathbb{L}},~{}~{}||\bm{z}||_{\mathbb{L}}=\sqrt{{\langle\bm{z},\bm{z}\rangle}_{\mathbb{L}}}.$ The logarithmic map is the inverse of this mapping and can be described as $\log_{\bm{x}}^{K}(\bm{y})=\frac{\cosh^{-1}(\beta)}{\sqrt{\beta^{2}-1}}\cdot(\bm{y}-\beta\bm{x}),\text{ with }\beta=-\frac{1}{K}\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}.$
3.2 Riemannian Optimization
On the Stability of Curvature Learning
@@ 3.2 Riemannian Optimization
In the hyperbolic learning library GeoOpt, Kochurov et al. (2020) introduce the curvature of hyperbolic space as a learnable parameter. However, no subsequent work has fully leveraged this feature, and our empirical testing reveals that this approach often leads to instability and performance decline. We attribute these issues to the simplistic implementation of the curvature updates, which neglects to update the hyperbolic operations after updating the curvature.
@@
In the current Riemannian optimization approach, Euclidean gradients and their momentums are projected onto the Riemannian manifold and then used to update the hyperbolic parameters. These operations inherently depend on the curvature of the space. However, when the curvature is adjusted during learning, the projections, gradients, and hyperbolic parameters can become inconsistent, as they mix the projection and update operations from the new curvature with vectors defined under the previous curvature.
@@
To illustrate this, Figure 1 shows two manifolds with different curvatures. It is evident that points defined on the first manifold do not lie on the second, and the tangent planes of the first manifold are not aligned with those of the second. Thus, using the new curvature operations on points from the old model is not mathematically sound practice. This inconsistency introduces instability, undermining the reliability of curvature learning and often leading to invalid values during the parameter updates. To address this, we propose a method that tracks both the old and new curvature values throughout updates. Specifically, all updates are applied under the old curvature, after which the hyperbolic vectors are mapped from the old manifold to the new one.
@@
While GeoOpt does not address the optimization challenges involved in learning manifold curvature, it does implement an ”N-stabilize step”. This step aims to reduce instability by recalculating the time component of the hyperbolic parameters every N-steps to ensure they remain on the manifold. However, this approach fails to prevent invalid parameter values before the update steps and suffers from high inaccuracies due to the mathematically flawed parameter optimization.
@@
Moreover, the time re-calculation projection used in the stabilize step alters the relative magnitudes of the hyperbolic vectors and gradients, and even the directions of the momentums, which can degrade performance and lead to training instability. Instead, we recommend projecting points onto the tangent space at the old origin using the logarithmic map and then projecting back after the curvature update. Gradients and their momentum can be similarly projected or parallel transported to the tangent space. Parallel transport is chosen here because it preserves the vectors’ directions which is important for the optimization procedure.
Additionally, when the curvature changes, the tangent space at the origin remains relatively stable, as shown in Figure 1. Moving onto the new manifold is then a simple exponential map and parallel transport back operation. This preserves the vector norms and their hyperbolic distances to the origin since $D(\exp^{K_{1}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{1}}=||y||=D(\exp^{K_{2}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{2}}$ where $\bm{y}\in\mathcal{T}_{\overline{\bm{0}}}\mathcal{M}$. No changes are needed for the tangent vector during the move between tangent spaces of different curvatures since these spaces are parallel, making the transition a simple translation along the time axis.
(a) At a random point
(b) At the origin
Figure 1: Tangent planes of a hyperboloid with curvature -1 relative to another hyperboloid with curvature -0.7. Tangential properties between manifolds are better respected at the origin where tangents remain parallel.
@@
Riemannian AdamW Optimizer
Recent works, particularly with transformers, commonly use the AdamW optimizer introduced by Loshchilov & Hutter (2019). As of current, there is no established Riemannian variant of this optimizer. We attempt to derive AdamW for the Lorentz manifold and argue a similar approach could be extended to the Poincaré ball. The main difference between AdamW and Adam is the direct weight regularization. This is more difficult to perform in the Lorentz space given the lack of an intuitive subtraction operation on the manifold. To resolve this, we reframe the parameter regularization in the original AdamW as a weighted centroid with the origin $O$
$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda O$

### v3

# Robust Hyperbolic Learning with Curvature-Aware Optimization

@@ Abstract
Hyperbolic deep learning has become a growing research direction in computer vision due to the unique properties afforded by the alternate embedding space. The negative curvature and exponentially growing distance metric provide a natural framework for capturing hierarchical relationships between datapoints and allowing for finer separability between their embeddings. However, current hyperbolic learning approaches are still prone to overfitting, computationally expensive, and prone to instability, especially when attempting to learn the manifold curvature to adapt to tasks and different datasets. To address these issues, our paper presents a derivation for Riemannian AdamW that helps increase hyperbolic generalization ability. For improved stability, we introduce a novel fine-tunable hyperbolic scaling approach to constrain hyperbolic embeddings and reduce approximation errors. Using this along with our curvature-aware learning schema for Lorentzian Optimizers enables the combination of curvature and non-trivialized hyperbolic parameter learning. Our approach demonstrates consistent performance improvements across Computer Vision, EEG classification, and hierarchical metric learning tasks achieving state-of-the-art results in two domains and drastically reducing runtime.
Machine Learning, ICML
@@ 1 Introduction
Recently, hyperbolic manifolds have gained attention in deep learning for their ability to model hierarchical and tree-like data structures efficiently. Unlike Euclidean space, hyperbolic space has negative curvature, allowing it to represent exponentially growing distances. This makes these manifolds ideal for tasks like natural language processing, graph representation, and recommendation systems (Peng et al., 2022). By embedding data into hyperbolic space, models can capture complex relationships more effectively, often with fewer parameters (Ganea et al., 2018), leading to improved performance and interpretability. Hyperbolic geometry has also found applications in computer vision, where its ability to better separate embeddings in high-dimensional spaces has shown promise in improving tasks like image classification and segmentation(Bdeir et al., 2024; van Spengler et al., 2023; Khrulkov et al., 2020; Atigh et al., 2022), few-shot learning(Liu et al., 2020), and feature representation (Yang et al., 2023).
Two main derivations of hyperbolic space have emerged as learning frameworks: the hyperboloid or the Lorentz space ($\mathbb{L}$) and the Poincaré manifold ($\mathbb{P}$). Typically, the hyperboloid, or Lorentz space, offers better operational stability, as demonstrated by Mishne et al. (2022) but lacks clear definitions for basic vector operations such as addition and subtraction. To bridge this gap, recent research has focused on defining Lorentzian versions of common deep learning operations, such as the feed-forward layer (Chen et al., 2022; Dai et al., 2021; Ganea et al., 2018), convolutional layer (Chen et al., 2022; Qu & Zou, 2023; Dai et al., 2021), and MLRs (Bdeir et al., 2024).
However, the use of these hyperbolic operations comes with challenges. Computations in hyperbolic space are more complex and expensive due to the increased mathematical complexity and the lack of optimized CUDA implementations, which can slow down training and drastically increase memory requirements. Additionally, optimization on hyperbolic manifolds can be unstable, especially in low floating-point precision environments. This has required research to rely on parameter clamping techniques, which can cause non-smooth gradient updates, to remain within the accurate representation radius for manifolds Mishne et al. (2022); Guo et al. (2022). Otherwise, performance either suffers or training becomes impossible due to the precision errors causing mathematically undefined operations. The instability is only exacerbated when we incorporate the negative curvature as a learned parameter since it directly affects the representational radius and requires the projection of vectors between these different curvatured manifolds. Finally, in data-scarce scenarios, the higher representational capacity makes the hyperbolic spaces more prone to overfitting which heavily impacts their generalization ability.
This work addresses key challenges in the Lorentz model. To combat overfitting, we introduce an AdamW-based optimizer that enhances regularization and generalization in hyperbolic learning. For learning instability, we propose an optimization framework that stabilizes curvature parameter learning. We also present a smoother scaling function to replace weight clipping, ensuring hyperbolic vectors remain within the representational radius. To tackle computational complexity, we introduce an implementation trick that leverages efficient CUDA-based convolutions for hyperbolic learning. Improved stability not only boosts model performance but also enables lower-precision training, enhancing efficiency. Our contributions are four-fold:
@@
    We propose a formulation for Riemannian AdamW and an alternative schema for Riemannian optimizers that stabilize curvature learning for hyperbolic parameters.
@@
    We present HCNN+, HyperMAtt, and LHEIR, an application of our proposed optimization scheme for the domains of classical image classification, EEG classification, and hierarchical metric learning respectively.
@@
    We empirically show the effectiveness of our proposed methods in three domains, hierarchical metric learning, Image, and EEG classification, where we achieve state-of-the-art performance in two domains with a significant computational speed-up.
Our code is publicly available at [redacted for double-blind submission].
@@ 2 Related Work
Hyperbolic Embeddings in Deep Learning
@@ 3 Methodology
3.1 The Lorentz Space
The hyperbolic space is a Riemannian manifold with a constant negative sectional curvature $c<0$. There are many conformal models of hyperbolic space but we focus our work on the hyperboloid, or Lorentz manifold. The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}=[x_{t},\bm{x_{s}}]\in\mathbb{R}^{n+1}~{}|~{}\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=\frac{-1}{K},~{}x_{t}>0\}$ where $\frac{-1}{K}=c$ and the Lorentzian inner product as Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}.$ It then follows that ${}\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}<\frac{-1}{K}$ which is an important characteristic to note for the stability issues presented later on. The formulation $\mathcal{L}^{n}$ then presents the Lorentz manifold as the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}=[\sqrt{K},0,\cdots,0]^{T}$.
We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector, as the space dimension $\bm{x_{s}}$. Below are the basic operations presented by the manifold with more complex operations provided in Appendix A.
@@ 3.1 The Lorentz Space
Distance in hyperbolic space is the magnitude of the geodesic forming the shortest path between two points. Let $\bm{x},\bm{y}\in\mathbb{L}^{n}_{K}$, the distance between them is given by
$d_{\mathbb{L}}(\bm{x},\bm{y})=\sqrt{K}\cosh^{-1}(\frac{-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}{K})$ (1)
We also define the square distance by Law et al. (2019) as $d^{2}_{\mathbb{L}}(\bm{x},\bm{y})=||\bm{x}-\bm{y}||_{\mathbb{L}}^{2}=-2K-2\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$.
@@
Since the Lorentz space is a Riemannian manifold, it is locally Euclidean. This can best be described through the tangent space $\mathbb{T}_{\bm{x}}\mathcal{M}$, a first-order approximation of the manifold at a given point $\bm{x}$. The exponential map, $\exp^{K}_{\bm{x}}\left(\bm{z}\right):\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}\rightarrow\mathbb{L}^{n}_{K}$ is then the operation that maps a tangent vector $\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}$ onto the manifold through $\exp_{\bm{x}}^{K}(\bm{z})=\cosh(\alpha)\bm{x}+\sinh(\alpha)\frac{\bm{z}}{\alpha},\text{ with }\alpha=\sqrt{1/K}||\bm{z}||_{\mathbb{L}},~{}~{}||\bm{z}||_{\mathbb{L}}=\sqrt{{\langle\bm{z},\bm{z}\rangle}_{\mathbb{L}}}.$ The logarithmic map is the inverse of this mapping and can be described as $\log_{\bm{x}}^{K}(\bm{y})=\frac{\cosh^{-1}(\beta)}{\sqrt{\beta^{2}-1}}\cdot(\bm{y}-\beta\bm{x}),\text{ with }\beta=-\frac{1}{K}\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}.$
3.2 Riemannian Optimization Schema
Background
@@ 3.2 Riemannian Optimization Schema
In the GeoOpt hyperbolic learning library, Kochurov et al. (2020) introduces the curvature $c$ of hyperbolic space as a learnable parameter. However, no subsequent work has fully taken advantage of this feature despite the authors presenting ready implementations of hyperbolic operations that account for this curvature. Our empirical testing also reveals that naively learning this parameter often leads to instability and performance decline. We attribute these issues to the simplistic curvature updates in Riemannian optimizers, which neglect to adjust hyperbolic operations, weights, and gradients after updating the curvature.
@@
In the current Riemannian optimization approach, Euclidean gradients and momentums are projected onto the manifold to update hyperbolic parameters. However, since optimization depends on curvature, adjusting it during learning makes previous projections, gradients, and parameters inconsistent with the new geometry. Essentially, the current schema applies updates using the new curvature on vectors defined by the old curvature, leading to mismatches.
@@
To illustrate this, LABEL:fig:curve_inconsistencies shows two manifolds with different curvatures. It is evident that points defined on the first manifold do not lie on the second, and the tangent planes of the first manifold are not aligned with those of the second. Thus, using the new curvature operations on points from the old model is not mathematically sound practice. This inconsistency introduces instability, undermining the reliability of curvature learning and often leading to undefined values during the parameter updates.
One example of this is the distance function Equation 1 which requires the bound ${}\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}<\frac{-1}{K}$ to hold. If the new curvature parametrization at time t $K_{t}>K_{t-1}$ then the input into the inverse hyperbolic cosine has a lower bound less than 1, leaving room for the operation to be mathematically undefined.
@@
While GeoOpt does not address this particular optimization challenge, it does implement an ”N-stabilize step”. This step aims to reduce instability by recalculating the time component of the hyperbolic parameters every N-steps to ensure they remain on the manifold. However, this approach is done after the update step and does not prevent invalid parameter values during optimization.
Curvature Aware Optimization
To address this, we propose a method that tracks both the old and new curvature parameter values $\bm{\theta}_{K}$ throughout updates. Specifically, we propose a simple approach where all updates are first applied under the old curvature, then the new curvature is derived and finally, the hyperbolic vectors are mapped from the old manifold to the new one.
@@
The previously proposed mapping technique in the N-stabilize step alters the relative magnitudes of the hyperbolic parameters $\bm{\theta}_{\mathbb{L}}$ and gradients $\mathcal{G}$, as well as the directions of the momentums, which can degrade performance and lead to training instability. Instead, we recommend projecting points onto the tangent space at the origin of the old curvature $\overline{\bm{0}}_{t-1}$ using the logarithmic map and then projecting back after the curvature update. Gradients and their momentum can be similarly projected or parallel transported Appendix A) to the tangent space. The entire optimization schema can then be seen in Algorithm 1.
Here, we choose to apply parallel transport, or any similar approximation of it, because it preserves the gradient vectors’ directions which is important for the optimization procedure. We also choose to use the tangent space at the origin as the mapping intermediate since it remains relatively stable when the curvature is updated, as shown in LABEL:fig:curve_inconsistencies. Finally, we prove that the vector norms and their hyperbolic distances to the origin are preserved since $D(\exp^{K_{1}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{1}}=||y||=D(\exp^{K_{2}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{2}}$ where $\bm{y}\in\mathcal{T}_{\overline{\bm{0}}}\mathcal{M}$.
@@
3.3 Riemannian AdamW Optimizer
Background
The AdamW optimizer was first introduced by Loshchilov & Hutter (2019) and relies on an improved application of the L2-regularization factor in the base Adam optimizer. L2-regularization works on the principle that networks with smaller weights tend to have better generalization performance than equal networks with higher weight values. Adam applied the L2-regularization during the gradient update step by incorporating it in the loss. However, Loshchilov & Hutter (2019) argue that this is inconsistent since the regularization effect is reduced by the magnitude of the gradient norms. Instead, they apply the weight decay directly during the parameter update step. AdamW is then shown to generalize better and lead to better convergence and has become a popular choice for many vision tasks.
Given the above, we believe AdamW is significant for hyperbolic learning. Hyperbolic spaces, with their higher representational capacity, are prone to overfitting, especially under data scarcity during training (Gao et al., 2022). Thus, AdamW’s enhanced L2-regularization could improve hyperbolic models.
Riemannian AdamW
In the following, we derive AdamW for the Lorentz manifold and suggest its extension to the Poincaré ball. The key difference between AdamW and Riemannian Adam lies in direct weight regularization, which is challenging in Lorentz space due to the lack of an intuitive subtraction operation. To address this, we re-frame parameter regularization in AdamW as a weighted centroid with the origin $\overline{\bm{0}}$
$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda\overline{\bm{0}}$